Race One Motors is an Indonesian car manufacturer. At its largest manufacturing facility, in Jakarta, the company produces subcomponents at a rate of 285 per day, and it uses these subcomponents at a rate of 12,300 per year (of 250 working days). Holding costs are $2 per item per year, and ordering costs are $32 per order
ROUND TO TWO DECIMAL POINTS FOR ALL ANSWERS
A.What is the economic production quantity ?

B. How many production runs per year will be made?

C. What will be the maximum inventory level?

D. What percentage of time will the facility be producing components?

E. What is the annual cost of ordering and holding inventory?
Answer #1

Given, daily production rate, p= 285 subcomponents

Annual usage rate/ demand, D= 12300 subcomponents

Number of days working in a year= 250

Hence daily usage rate, d= 12300/250= 49.2 subcomponents

Holding cost per unit per year, H= $2

Ordering cost, S= $32

A. Economic production quantity, Q =

=

= 689.73 subcomponents

B. Number of production runs per year= D/Q = 12300/689.73= 17.83

C. Maximum inventory level= Q(1-d/p) = 689.73(1-(49.2/285)) = 570.66

D. The percentage of time will the facility be producing components= (length of production run*number of production runs per year)/ total number of working days in an year

Length of production run= Q/p = 689.73/285 = 2.42

Hence the percentage of time will the facility be producing components= (2.42*17.83)/250 = 17.26%

E. Annual cost of ordering= (D/Q)S = (12300/689.73)*32 = $570.66

Annual cost of holding the inventory= QH/2(1-(d/p))= 689.73*2/2(1-49.2/285)= $570.66